OpenStudy (anonymous):

suppose you wanted to place 1,000,000 pennies in large coin banks. if you put 50,000 pennies in each bank, in which bank would you find the 678,230th penny?

OpenStudy (anonymous):

678 230 / 50 000 = about 13,56. It's over 13 and less then 14, so then its the 14th bank. (50 000*13=650 000 and 50*14=700 000, therefor it's in the 14th bank)

OpenStudy (anonymous):

thank u sjano

OpenStudy (anonymous):

No problem. Just remember to close the question :P